40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 14 21:37:37 2012, yuta, Update, Green Locking, succeeded in 1FSR mode scan YarmScan20120614.png
    Reply  Fri Jun 15 00:50:54 2012, yuta, Update, Green Locking, scanned Y arm for 5FSR YarmScan20120614_2.png
       Reply  Fri Jun 15 01:53:05 2012, Koji, Update, Green Locking, scanned Y arm for 5FSR 
          Reply  Sat Jun 16 12:03:41 2012, Zach, Update, Green Locking, scanned Y arm for 5FSR 
             Reply  Sat Jun 16 13:01:17 2012, yuta, Update, Green Locking, scanned Y arm for 5FSR YarmScanDV.png
       Reply  Sat Jun 16 18:17:00 2012, yuta, Update, Green Locking, Y arm length using 5FSR scan ALS-BEATY_COARSE_I_IN1_DQ.pngCalibYarmScan20120614_2.png
          Reply  Sat Jun 16 18:51:44 2012, Koji, Update, Green Locking, Y arm length using 5FSR scan 
             Reply  Sat Jun 16 19:32:11 2012, yuta, Update, Green Locking, Y arm length using 5FSR scan 
Message ID: 6818     Entry time: Thu Jun 14 21:37:37 2012     Reply to this: 6819
Author: yuta 
Type: Update 
Category: Green Locking 
Subject: succeeded in 1FSR mode scan 

[Jenne, Yuta]

We couldn't scan the Y arm for 1FSR last night because the ALS servo breaks while sweeping.
We thought this might be from the amplitude fluctuation of the beat signal. The amplitude of the beat signal goes into the beatbox was about -5 dBm, which is not so enough for the beatbox to get good LO. So, we put an amplifier (and attenuators) and the amplitude became +1 dBm. The range beatbox can handle is about -3 dBm to +3 dBm, according to our calculation.

This increased stability of the lock, and we could scan the arm for 1FSR. Below is the plot of scanned ALS error signal (blue), Y arm IR PDH signal (green) and TRY (red).

YarmScan20120614.png

For each slope, we can see two TEM00 peaks, some higer order modes(may be 01, 02, 02) and sidebands (large 11MHz, small 55MHz?).

We couldn't scan for more. This is still a mystery.

Also, we need to reduce residual Y arm length fluctuation more because we get funny TRY peak shape.

Scan speed:
  For C1:ALS-BEATY_COARSE_I_IN1, 1 count stands for 0.21 nm(see elog #6817). We sweeped 4000 peak to peak in 50 sec. So, the scan speed is about 17 nm/sec.
  This means it takes about 0.06 sec to cross resonant peak.
  Cavity build up time is about 2LF/(pi*c) ~ 40 usec. So, the scan is quasi-static enough.
  Characteristic time scale for the Y end temperature control is about 10 sec, so Y end frequency is following the Y arm length change with temperature control.

  Currently, sampling frequency of DQ channels are 2048 Hz. This means we have 100 points in a TRY peak. I think this is enough to get a peak height.

Next step:
  - Reduce RMS. We are trying to use a whitening filter.
  - Find why we can't scan more. Why??
  - ETMY coil gains may have some unbalance. We need to check
  - Characterize Y end green frequency control. Koji and I changed them last week (see elog #6776).
  - Calculate positions of RF SBs and HOMs and compare with this result.

ELOG V3.1.3-